# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 5.1 Setup
I also calculate the mean plug-in estimated asymptotic variance and the oracle scaled variance (similar to $\sigma_{\dagger,a}^{2}$ and $\sigma_{\#,a}^{2}$ when estimating $\psi_{*,a}$) for $Q_{\#,a}$ being the mean of $\hat{Q}_{k,a}$ or $\tilde{Q}_{a}$, that is, satisfying Part 2 of Condition 8 or 11. For each AIPW estimator, the mean estimated asymptotic variance is approximated by the Monte Carlo mean over simulation runs. It is more challenging to calculate $\sigma_{\#,a}^{2}$. I describe its calculation in Section S3 of the Supplementary Material.
@@ 5.2 Results
Fig. 2 and 3 present the sampling distributions of the ATE estimators and the corresponding estimated asymptotic variance, respectively. Fig. 4 presents the QQ-plot of the AIPW estimators. As expected, all estimators appear approximately unbiased; IPW has the largest variance because predictive covariates are not used; due to less bias in estimating $Q_{*,a}$, SL and CVSL have smaller variance than ANCOVA, misSL, and CVmisSL.
@@
Fig. 5 presents the CI coverage. SL severely undercovers, and a plausible reason is the underestimated variance. All other CIs appear to have approximately correct coverage, particularly when $n\geq 200$, aligning with the results and discussions in Sections 3 and 4. Although sometimes CVSL has a coverage that is statistically significantly lower than 95%, the coverage is arguably practically close to 95%, particularly when $n\geq 200$. Interestingly, with more covariates, CVSL appears to have higher coverage. A possible reason is that, with more covariates, $\hat{Q}_{k,a}$ has more variability, so that the estimated variance has a positive bias by (3), which might remedy undercoverage. The potential trade-off between efficiency and Wald-CI coverage shows up for the non-cross-fit procedure in the simulation: misSL is less efficient than SL but has better Wald-CI coverage. This trade-off is less clear for the cross-fit procedure in the simulation. A potential reason is the missing empirical process term.
@@ 6 Discussion
The theoretical results have room for improvement. First, the bounds for empirical processes might be overly conservative. It might be possible to substantially reduce those bounds with some additional high-level conditions. Second, there remains a theoretical gap regarding how the bias of the estimated variance impacts the CI coverage. Although the simulation and my argument in Section 4 indicate that cross-fitting may increase CI coverage and non-cross-fitting may decrease CI coverage, this bias is not a first-order term in my Berry-Esseen-type bounds in Theorems 1 and 3. Third, I derived the bounds based on empirical process theory and tail bounds. For simpler estimators such as sample mean, methods based on characteristic functions yield sharper characterization (without log factors) under other mild conditions (e.g., the Berry-Essen Theorem). I did not adopt this approach to accommodate the black-box nuisance function estimators. With additional knowledge on the nuisance estimators, it is possible to use characteristic functions to achieve a more accurate characterization of the CI coverage convergence rate.
@@
Nevertheless, I provide new theoretical justifications for adopting cross-fitting when using an AIPW estimator with a flexible outcome model estimator to estimate the ATE in an RCT. On the one hand, my upper bound of the convergence rate of CI coverage to nominal coverage for cross-fitting vanishes faster than the non-cross-fit version. On the other hand, cross-fitting overestimates the variance, providing additional protection against undercoverage. The non-asymptotic results in Section S2 of the Supplementary Material may also enable asymptotic analysis in high-dimensional settings. The results of this paper lay the foundation for future research on the coverage of CIs based on non- and semiparametric estimators, and potentially for providing practical guidance on using such procedures.
S1 Additional conditions
Condition S1 (Additional subgaussian condition for $\hat{Q}_{k,a}$).
For all $k\in\{1,\ldots,K\}$ and $a\in\{0,1\}$,
$\frac{P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]}{M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}/\tau_{\pi}^{2}}$
is subgaussian with a variance proxy $\hat{s}_{3}^{2}$ (namely a subgaussian parameter $\hat{s}_{3}>0$).
Condition S2 (Additional subgaussian condition for $\tilde{Q}_{a}$).
For all $a\in\{0,1\}$,
$\frac{P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]}{M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}/\tau_{\pi}^{2}}$
is subgaussian with a variance proxy $\tilde{s}_{2}^{2}$ (namely a subgaussian parameter $\tilde{s}_{2}>0$).
These two conditions are similar. I illustrate Condition S1 next. As shown in Lemma S6, the variance of the numerator $P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\hat{Q}_{k,a})^{2}]$ is upper bounded by the denominator $M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}/\tau_{\pi}^{2}$ (up to an absolute constant 16) in the this condition. In addition, because of the integration with respect to $P_{*}$ and the boundedness of $\hat{Q}_{k,a}$ in Condition 4, the random variable in the numerator is essentially a weighted average of the difference $\hat{Q}_{k,a}(X)-Q_{\#,a}(X)$ (see Lemma S6). Thus, this condition is essentially a restriction on the tail probability of an overall difference between $\hat{Q}_{k,a}$ and $Q_{\#,a}$.
S2 Additional theoretical results
@@
Suppose that $\sigma_{\dagger,a}>0$, i.e., $\Pr(\hat{\sigma}_{a}>0)>0$. Under Conditions 15, for any constant $q>0$, if $\psi_{*,a}=0$ or $z_{\alpha}=0$ (i.e., $\alpha=0.5$) or $2(\sqrt{2}+1)^{2}q\bar{\sigma}_{a}^{2}\log n/|I_{k}|\leq\psi_{*,a}^{2}$ for all $k=1,\ldots,K$, then
$\left|\Pr(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$\leq{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+{\mathcal{C}}\sum_{k=1}^{K}\Bigg[\hat{\mathcal{R}}_{k}+\frac{\hat{S}_{k}^{2}}{\sigma_{\#,a}^{4}}+\left(\frac{\bar{m}}{\underline{m}^{3/2}}+{\mathbbm{1}}(z_{\alpha}\neq 0)\frac{\bar{\rho}}{\sigma_{*,a}^{3/2}}\right)\frac{1}{\sqrt{|I_{k}|}}+\frac{|z_{\alpha}|\sqrt{2q|I_{k}|\bar{m}\log n}}{n\sigma_{\#,a}\sigma_{\dagger,a}}+n^{-q}$
$\qquad+\frac{|I_{k}||z_{\alpha}|}{n\sigma_{\#,a}\sigma_{\dagger,a}}\left\{|\psi_{*,a}|\sqrt{q\frac{\bar{\sigma}_{a}^{2}}{|I_{k}|}\log n}+{\mathbbm{1}}(\psi_{*,a}=0)q\frac{\bar{\sigma}_{a}^{2}}{|I_{k}|}\log n+\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2}\right\}\Bigg].$ (S1)
@@
$\hat{\mathcal{R}}_{k}$ $:=\left[\left\{\frac{|I_{k}|}{n\sigma_{\#,a}^{2}}\frac{1-\tau_{\pi}}{\tau_{\pi}}\right\}^{1/3}+\left\{\frac{|I_{k}|^{2}M^{2}z_{\alpha}^{2}}{n^{2}\sigma_{\#,a}^{2}\sigma_{\dagger,a}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}\right\}^{1/3}\right]\left\{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3},$
$\hat{S}_{k}$ $:=\frac{|I_{k}|}{n}\Bigg\{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+2\frac{1-\tau_{\pi}}{\tau_{\pi}}\|Q_{\#,a}-Q_{*,a}\|_{2}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}$
$\qquad\qquad+\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2}\Bigg\}.$
Moreover, under Condition 8, $\hat{S}_{k}$ can be replaced by
$\hat{S}_{k}^{\prime}:=\frac{|I_{k}|}{n}\Bigg\{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}+\frac{1}{|I_{k}|}\left\{\sigma_{\#,a}+\sqrt{\frac{1-\tau_{\pi}}{\tau_{\pi}}{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}\right\}^{2}\Bigg\};$
under subgaussian conditions 7 and S1, $\mathcal{R}_{k}$ can be replaced by
$\hat{\mathcal{R}}_{k}^{\prime}$ $:=\frac{\hat{s}_{1}}{\sigma_{\#,a}}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+\hat{s}_{2}\sqrt{q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}\right]\sqrt{q\frac{|I_{k}|}{n}\log n}$
$\qquad+\frac{|I_{k}|\hat{s}_{3}|z_{\alpha}|M(1-\tau_{\pi})}{n\sigma_{\#,a}\sigma_{\dagger,a}\tau_{\pi}}\sqrt{q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}.$
The next theorem is a non-asymptotic version of Theorem 2 and Part 1 of Proposition 1. Theorem 2 follows by noting that ${\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}\leq\sqrt{{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}$ and ${\mathrm{Var}}(\hat{\psi}_{k,a})={\mathrm{O}}(n^{-1})$ by Lemma S8 under the setting of Theorem 2. Part 1 of Proposition 1 follows by Lemma S8.
@@
Consider the cross-fit procedure. Under Conditions 1 and 3, $|\sigma_{\dagger,a}-\sigma_{\#,a}|\leq\sum_{k=1}^{K}\hat{S}_{k}/\sigma_{\#,a}$. Additionally under Condition 8,
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=\sum_{k=1}^{K}\frac{|I_{k}|}{n}\left\{{\mathbb{E}}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(\hat{Q}_{k,a}-Q_{\#,a})^{2}\right]-{\mathrm{Var}}(\hat{\psi}_{k,a})\right\}$ (S2)
and $|\sigma_{\dagger,a}-\sigma_{\#,a}|\leq\sum_{k=1}^{K}\hat{S}_{k}^{\prime}/\sigma_{\#,a}$.
@@
Suppose that $\sigma_{\dagger,a}>0$, i.e., $\Pr(\tilde{\sigma}_{a}>0)>0$. Under Conditions 15 and 9, for any constants $\delta,\delta^{\prime}>0$ and $q\geq 1$,
$\left|\Pr(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$\leq{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}\sqrt{n}}+{\mathcal{C}}{\mathbbm{1}}(z_{\alpha}\neq 0)\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}\sqrt{n}}+{\mathcal{C}}n^{-q}+\tilde{\mathcal{R}}(\delta,\delta^{\prime})+$
$\quad+\frac{{\mathcal{C}}}{\sigma_{\#,a}^{4}}\Bigg[\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2\delta^{\prime},{\mathcal{F}},M)}{n}+\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{4}(2\delta^{\prime},{\mathcal{F}},M)}{(\delta^{\prime})^{4}n^{2}}+\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{2}$
$\qquad\qquad+\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}\|Q_{\#,a}-Q_{*,a}\|_{2}^{2}({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}+\frac{\sigma_{\#,a}^{4}+(F^{H_{a}})^{4}J^{4}(2,{\mathcal{F}},M)}{n^{2}}\Bigg]$
$\quad+{\mathcal{C}}\frac{|z_{\alpha}|}{\sigma_{\#,a}\sigma_{\dagger,a}}\Bigg[\frac{\sigma_{\#,a}^{2}q\log n}{n}+(|\psi_{*,a}|\sigma_{\#,a}+\varsigma_{\#,a})\sqrt{\frac{q\log n}{n}}+\frac{F^{H_{a}}\sigma_{\#,a}\sqrt{q\log n}}{n}$
$\qquad\qquad+\frac{(F^{H_{a}})^{2}\{J^{2}(2,{\mathcal{F}},M)+q^{2}(\log n)^{2}\}}{n}+\frac{|\psi_{*,a}|F^{H_{a}}\{J(2,{\mathcal{F}},M)+q\log n\}}{\sqrt{n}}$
$\qquad\qquad+\left(F^{H_{a}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)J(2\delta,{\mathcal{F}},M)+\frac{(F^{H_{a}}+F^{{\mathcal{T}}_{a}^{2}}n^{-1/2})J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}}$
$\qquad\qquad+q\left(F^{H_{a}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)(\delta+n^{-1/2})\log n\Bigg].$ (S3)

### v2

@@ 5.1 Setup
I also calculate the mean plug-in estimated asymptotic variance and the oracle scaled variance (similar to $\sigma_{\dagger,a}^{2}$ and $\sigma_{\#,a}^{2}$ when estimating $\psi_{*,a}$) for $Q_{\#,a}$ being the mean of $\hat{Q}_{k,a}$ or $\tilde{Q}_{a}$, that is, satisfying Part 2 of Condition 9 or 13. For each AIPW estimator, the mean estimated asymptotic variance is approximated by the Monte Carlo mean over simulation runs. It is more challenging to calculate $\sigma_{\#,a}^{2}$. I describe its calculation in Section S4 of the Supplementary Material.
@@ 5.2 Results
Figures 1 and 2 present the sampling distributions of the ATE estimators and the corresponding estimated asymptotic variance, respectively. Figure 3 presents the QQ-plot of the AIPW estimators. As expected, all estimators appear approximately unbiased; IPW has the largest variance because predictive covariates are not used; due to less bias in estimating $Q_{*,a}$, SL and CVSL have smaller variance than ANCOVA, misSL, and CVmisSL.
@@
Figure 4 presents the CI coverage. SL severely undercovers, and a plausible reason is the underestimated variance. All other CIs appear to have approximately correct coverage, particularly when $n\geq 200$, aligning with the results and discussions in Sections 3 and 4. Although sometimes CVSL has a coverage that is statistically significantly lower than 95%, the coverage is arguably practically close to 95%, particularly when $n\geq 200$. Interestingly, with more covariates, CVSL appears to have higher coverage. A possible reason is that, with more covariates, $\hat{Q}_{k,a}$ has more variability, so that the estimated variance has a positive bias by (3), which might remedy undercoverage. The potential trade-off between efficiency and Wald-CI coverage shows up for the non-cross-fit procedure in the simulation: misSL is less efficient than SL but has better Wald-CI coverage. This trade-off is less clear for the cross-fit procedure in the simulation. A potential reason is the missing empirical process term.
@@ 6 Discussion
The theoretical results might have room for improvement. First, the bounds for empirical processes might be overly conservative. For example, it might be possible to reduce those bounds with some additional high-level conditions. Second, there remains a theoretical gap regarding how the bias of the estimated variance impacts the CI coverage. Although the simulation and my argument in Section 4 indicate that cross-fitting may increase CI coverage and non-cross-fitting may decrease CI coverage, this bias is not a first-order term in my Berry-Esseen-type bounds in Theorems 1 and 3. Third, I derived the bounds based on empirical process theory and tail bounds. For simpler estimators such as the sample mean, methods based on characteristic functions yield sharper characterization (without log factors) under other mild conditions (e.g., the Berry-Essen Theorem). I did not adopt this approach to accommodate the black-box nuisance function estimators. With additional knowledge on the nuisance estimators, it is possible to use characteristic functions to achieve a more accurate characterization of the CI coverage convergence rate. Finally, for the cross-fitting procedure, although I treat the number of folds $K$ as a variable, as I discuss in Section S3 in the Supplementary Material, my bound might have room for improvement in terms of $K$. A theoretically sound recommendation of $K$ is an open question.
@@
Nevertheless, I provide new theoretical justifications for adopting cross-fitting when using an AIPW estimator with a flexible outcome model estimator to estimate the ATE in an RCT. On the one hand, my upper bound of the convergence rate of CI coverage to nominal coverage for cross-fitting vanishes faster than the non-cross-fit version. On the other hand, cross-fitting overestimates the variance, providing additional protection against undercoverage. The non-asymptotic results in Section S1 of the Supplementary Material may also enable asymptotic analysis in high-dimensional settings. The bounds under the observational setting are an interesting future direction, but they might require substantively different analyses, although the analytic tools in the paper might help. The results of this paper lay the foundation for future research on the coverage of CIs based on non- and semi-parametric estimators, and potentially for providing practical guidance on using such procedures.
S1 Additional theoretical results
In the following non-asymptotic results, the terms concerning the rates in the corresponding asymptotic scenario (Theorems 14), i.e., the terms depending on $(n,K,\hat{Q}_{k,a},\delta)$ in the upper bounds, are colored blue.
@@
Suppose that $\sigma_{\dagger,a}>0$, i.e., ${\mathrm{pr}}(\hat{\sigma}_{a}>0)>0$. Under Conditions 15, for any constant $q>0$, if $\psi_{*,a}=0$ or $z_{\alpha}=0$ (i.e., $\alpha=0.5$) or $|I_{k}|$ is sufficiently large so that $2(\sqrt{2}+1)^{2}q\bar{\sigma}_{a}^{2}{\color[rgb]{0,0,1}\log n/|I_{k}|}\leq\psi_{*,a}^{2}$ for all $k=1,\ldots,K$, then
$\left|{\mathrm{pr}}(\sqrt{n}(\hat{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\hat{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$\leq{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}{\color[rgb]{0,0,1}\sqrt{n}}}+{\mathcal{C}}{\color[rgb]{0,0,1}\sum_{k=1}^{K}}\Bigg[{\color[rgb]{0,0,1}\hat{\mathcal{R}}_{k}}+\frac{{\color[rgb]{0,0,1}\hat{S}_{k}^{2}}}{\sigma_{\#,a}^{4}}+\left(\frac{\bar{m}}{\underline{m}^{3/2}}+{\mathbbm{1}}(z_{\alpha}\neq 0)\frac{\bar{\rho}}{\sigma_{*,a}^{3/2}}\right){\color[rgb]{0,0,1}\frac{1}{\sqrt{|I_{k}|}}}+\frac{|z_{\alpha}|\sqrt{2q{\color[rgb]{0,0,1}|I_{k}|}\bar{m}{\color[rgb]{0,0,1}\log n}}}{{\color[rgb]{0,0,1}n}\sigma_{\#,a}\sigma_{\dagger,a}}+{\color[rgb]{0,0,1}n^{-q}}$
[…]
@@
$\hat{\mathcal{R}}_{k}$ $:=\left[\left\{\frac{{\color[rgb]{0,0,1}|I_{k}|}}{{\color[rgb]{0,0,1}n}\sigma_{\#,a}^{2}}\frac{1-\tau_{\pi}}{\tau_{\pi}}\right\}^{1/3}+\left\{\frac{{\color[rgb]{0,0,1}|I_{k}|^{2}}M^{2}z_{\alpha}^{2}}{{\color[rgb]{0,0,1}n^{2}}\sigma_{\#,a}^{2}\sigma_{\dagger,a}^{2}}\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}\right\}^{1/3}\right]{\color[rgb]{0,0,1}\left\{{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\right\}^{1/3}},$
[…]
Moreover, under Condition 9, $\hat{S}_{k}$ can be replaced by
[…]
under approximate sub-Weibull conditions 7 and 8, $\mathcal{R}_{k}$ can be replaced by
$\hat{\mathcal{R}}_{k}^{\prime}$ $:=\frac{1}{\sigma_{\#,a}}\left\{2[(q+\hat{d}_{1}){\color[rgb]{0,0,1}\log n}]^{(\hat{q}_{1}+2)/(2\hat{q}_{1})}\hat{c}_{1}^{-1/\hat{q}_{1}}{\color[rgb]{0,0,1}\hat{r}_{1}(n)}\sqrt{\frac{\tau_{\pi}}{1-\tau_{\pi}}{\color[rgb]{0,0,1}\frac{|I_{k}|}{n}}}\right\}\vee\left\{\frac{8}{3}M\frac{1-\tau_{\pi}}{\tau_{\pi}}\frac{(q+\hat{d}_{1}){\color[rgb]{0,0,1}\log n}}{{\color[rgb]{0,0,1}\sqrt{n}}}\right\}$
$\qquad+\left[\frac{(q+\hat{d}_{2}){\color[rgb]{0,0,1}\log n}}{\hat{c}_{2}}\right]^{1/\hat{q}_{2}}\frac{{\color[rgb]{0,0,1}|I_{k}|}|z_{\alpha}|}{{\color[rgb]{0,0,1}n}\sigma_{\dagger,a}}{\color[rgb]{0,0,1}\hat{r}_{2}(n)}+\frac{\{2\hat{a}_{1}+{\mathbbm{1}}(z_{\alpha}\neq 0)\hat{a}_{2}\}{\color[rgb]{0,0,1}n^{-q}}+\{2\hat{b}_{1}+{\mathbbm{1}}(z_{\alpha}\neq 0)\hat{b}_{2}\}{\color[rgb]{0,0,1}\sqrt{\log n/n}}}{\sigma_{\#,a}}.$
The next theorem is a non-asymptotic version of Theorem 2 and Part 1 of Proposition 1. Theorem 2 follows by noting that ${E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}\leq\sqrt{{E}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}}$ and ${\mathrm{var}}(\hat{\psi}_{k,a})={\mathrm{O}}(n^{-1})$ by Lemma S8 under the setting of Theorem 2. Part 1 of Proposition 1 follows by Lemma S8.
@@
Consider the cross-fit procedure. Under Conditions 1 and 3, $|\sigma_{\dagger,a}-\sigma_{\#,a}|\leq{\color[rgb]{0,0,1}\sum_{k=1}^{K}}{\color[rgb]{0,0,1}\hat{S}_{k}}/\sigma_{\#,a}$. Additionally under Condition 9,
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}={\color[rgb]{0,0,1}\sum_{k=1}^{K}}{\color[rgb]{0,0,1}\frac{|I_{k}|}{n}}\left\{{E}\left[P_{*}\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{\color[rgb]{0,0,1}(\hat{Q}_{k,a}-Q_{\#,a})^{2}}\right]-{\color[rgb]{0,0,1}{\mathrm{var}}(\hat{\psi}_{k,a})}\right\}$ (S2)
and $|\sigma_{\dagger,a}-\sigma_{\#,a}|\leq\sum_{k=1}^{K}{\color[rgb]{0,0,1}\hat{S}_{k}^{\prime}}/\sigma_{\#,a}$.
@@
Suppose that $\sigma_{\dagger,a}>0$, i.e., ${\mathrm{pr}}(\tilde{\sigma}_{a}>0)>0$. Under Conditions 15 and 10, for any constants $\delta,\delta^{\prime}>0$ and $q\geq 1$,
$\left|{\mathrm{pr}}(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
$\leq{\mathcal{C}}\frac{\rho_{\#,a}}{\sigma_{\#,a}^{3}{\color[rgb]{0,0,1}\sqrt{n}}}+{\mathcal{C}}{\mathbbm{1}}(z_{\alpha}\neq 0)\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}{\color[rgb]{0,0,1}\sqrt{n}}}+{\mathcal{C}}{\color[rgb]{0,0,1}n^{-q}}+{\color[rgb]{0,0,1}\tilde{\mathcal{R}}(\delta,\delta^{\prime})}$
$\quad+\frac{{\mathcal{C}}}{\sigma_{\#,a}^{4}}\Bigg[\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{2}(2\delta^{\prime},{\mathcal{F}},M)}{{\color[rgb]{0,0,1}n}}+\frac{(F^{{\mathcal{T}}_{a}^{2}})^{2}J^{4}(2\delta^{\prime},{\mathcal{F}},M)}{(\delta^{\prime})^{4}{\color[rgb]{0,0,1}n^{2}}}+\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}{\color[rgb]{0,0,1}({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{2}}$
$\qquad\qquad+\left(\frac{1-\tau_{\pi}}{\tau_{\pi}}\right)^{2}\|Q_{\#,a}-Q_{*,a}\|_{2}^{2}{\color[rgb]{0,0,1}({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}+\frac{\sigma_{\#,a}^{4}+(F^{H_{a}})^{4}J^{4}(2,{\mathcal{F}},M)}{{\color[rgb]{0,0,1}n^{2}}}\Bigg]$
$\quad+{\mathcal{C}}\frac{|z_{\alpha}|}{\sigma_{\#,a}\sigma_{\dagger,a}}\Bigg[\frac{\sigma_{\#,a}^{2}q{\color[rgb]{0,0,1}\log n}}{{\color[rgb]{0,0,1}n}}+(|\psi_{*,a}|\sigma_{\#,a}+\varsigma_{\#,a})\sqrt{\frac{q{\color[rgb]{0,0,1}\log n}}{{\color[rgb]{0,0,1}n}}}+\frac{F^{H_{a}}\sigma_{\#,a}\sqrt{q{\color[rgb]{0,0,1}\log n}}}{{\color[rgb]{0,0,1}n}}$
$\qquad\qquad+\frac{(F^{H_{a}})^{2}\{J^{2}(2,{\mathcal{F}},M)+q^{2}{\color[rgb]{0,0,1}(\log n)^{2}}\}}{{\color[rgb]{0,0,1}n}}+\frac{|\psi_{*,a}|F^{H_{a}}\{J(2,{\mathcal{F}},M)+q{\color[rgb]{0,0,1}\log n}\}}{{\color[rgb]{0,0,1}\sqrt{n}}}$
$\qquad\qquad+\left(F^{H_{a}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{{\color[rgb]{0,0,1}\sqrt{n}}}\right){\color[rgb]{0,0,1}J(2\delta,{\mathcal{F}},M)}+\frac{(F^{H_{a}}+F^{{\mathcal{T}}_{a}^{2}}{\color[rgb]{0,0,1}n^{-1/2}}){\color[rgb]{0,0,1}J^{2}(2\delta,{\mathcal{F}},M)}}{{\color[rgb]{0,0,1}\delta^{2}\sqrt{n}}}$
$\qquad\qquad+q\left(F^{H_{a}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{{\color[rgb]{0,0,1}\sqrt{n}}}\right){\color[rgb]{0,0,1}\left(\delta+\frac{1}{\sqrt{n}}\right)\log n}\Bigg].$ (S3)